# AgenticSciML: Collaborative Multi-Agent Systems for Emergent Discovery in Scientific Machine Learning

### v1

@@ 1 Introduction
Scientific Machine Learning (SciML) integrates data-driven learning with numerical simulation and physical modeling to solve problems in scientific computing that are analytically intractable or computationally expensive. Recent developments such as physics-informed neural networks (PINNs) [raissi2019physics], neural operators [li2020fourier, lu2021learning], and domain-decomposition methods [jagtap2020conservative, meng2020ppinn] have enabled progress across fluid mechanics, inverse problems, materials modeling, biophysics, and geoscience. However, the construction of effective SciML models still requires substantial manual effort: practitioners must select architectures, enforce physical constraints, tune loss weightings, design sampling curricula, and determine proper optimization strategies. This iterative search is labor-intensive and problem-specific, forming a bottleneck for scalable scientific computing.
Recent LLM-based agent systems have motivated interest in automating portions of the SciML workflow. Notable examples include the Virtual Lab [boiko2023virtual], in which domain-specialized agents collaborate to plan experiments and interpret results, and agent-based frameworks for molecular and materials design [bran2023chemist, gupta2023agent]. These approaches demonstrate the utility of role-specialization and iterative refinement, but they focus primarily on managing experimental or model-selection workflows. In parallel, AutoML systems and neural architecture search (NAS) frameworks [liu2018darts, real2019regularized] optimize hyperparameters or network topologies, while evolutionary and program-synthesis approaches such as AlphaEvolve [alphadevolve2024] and symbolic regression systems [cranmer2020discovering, udrescu2020ai] aim to discover compact representations. However, these systems do not generally produce new SciML modeling strategies, such as novel PDE-constrained architectures, problem-specific loss formulations, or adaptive decomposition schemes. Instead, they typically search within a predetermined hypothesis space or rely on a fixed modeling paradigm.
@@
To address this need, we introduce AgenticSciML, a collaborative multi-agent framework for the emergent discovery of SciML modeling strategies. The framework coordinates multiple specialized agents—including proposers, critics, engineers, retrievers, and evaluators—that iteratively generate, analyze, and refine candidate SciML solutions. A persistent knowledge base stores prior solutions, ablation outcomes, error analyses, and textual reports. Retrieval agents surface relevant methodological precedents, allowing new strategies to build upon earlier ones. A structured debate process encourages agents to justify, challenge, and revise modeling decisions, while an ensemble-guided evolutionary search balances exploitation of promising directions with exploration of alternative approaches.
@@ 2 Methods
Our evolutionary multi-agent system operates in three phases (Figure 1): user input, data analysis and evaluation criteria creation, and solution evolution. A team of 10+ agents are deployed with specialized roles. Key steps are detailed below.
@@
The problem and requirement statements are to provide contexts. The evaluation criteria define what success looks like for this problem, such as a low mean squared error on a validation set or low PDE residuals. Datasets are optional, and the user can choose to provide none, some, or all data needed for training or validation. Agents will use user-provided data when available, but can also generate synthetic data if needed. The user only needs to write in minimal prompts in natural language, and the agents will handle all technical details. See Appendix A for examples.
@@
Before proceeding to solution generation, an evaluator agent formalizes an evaluation contract that all candidate solutions must satisfy. The goal is to ensures consistent evaluation across diverse solutions and produces a scalar score for ranking. Note that throughout this work, the score is usually defined as a loss to minimize such as mean squared error (MSE) or relative $L_{2}$ error, so lower is better.
@@
The root engineer has access to all the user inputs and evaluation guidelines. It produces a baseline solution, named solution_0. The solution is executed and evaluated autonomously. The results, including training logs and any generated plots, are analyzed by a multi-modal result analyst agent, which writes a detailed textual report (analysis_0.md) saved to the Analysis Base for future reference.
@@
Planning: This is the critical stage for scientific reasoning and innovation. First, a retriever agent critically evaluates the parent’s solution and results analysis to determine if any entries from a curated Knowledge Base of SciML techniques would be helpful. If so, it retrieves 0-1 entries to share with the proposer and critic agents. More information on the Knowledge Base can be found in Appendix B.
@@ 3 Results
    Discontinuous Function Approximation (A.1): Learning a piecewise oscillatory function with discontinuities from data using neural networks.
@@
    Solving Poisson Equation on L-shaped Domain with PINN (A.2): Using Physics-Informed Neural Networks (PINNs) to solve the Poisson equation on an L-shaped domain.
@@
    Solving Burger’s Equation using PINN (A.3): Using PINNs to solve the time-dependent nonlinear Burger’s equation with periodic boundary conditions and a challenging initial condition.
@@
    Antiderivative Operator Learning (A.4): Learning a neural operator that maps input functions to their antiderivatives, generalizing to unseen functions.
@@
    Multiple-input Operator Learning for Reaction-Diffusion Equation (A.5): Learning a neural operator that maps multiple input functions (diffusion coefficient and source term) to the solution of a reaction-diffusion equation over time.
@@
    Reconstruction of 2D Cylinder Wake Vorticity Field from Sparse and Noisy Observations (A.6): Using neural networks to reconstruct the vorticity field behind a cylinder from only 4 sensors and noisy measurement.
These problems are designed to encourage original thinking from the agents and be solved with novel SciML techniques. Detailed setups and results for each problem are provided in the Appendix. All problems use the same knowledge base and similar agent parameters. Below, we summarize key findings across all experiments.
@@ 3.2 Novel Solution Strategies Emerge Through Evolution
    Discontinuous Function Approximation (A.1.4): A Mixture-of-Experts architecture with specialized experts for different branches of the piecewise function, controlled by a learnable gating mechanism. The trainable gating parameter is inspired by the trainable parameter in adaptive activation functions [jagtap2020adaptive].
@@
    Poisson Equation on L-shaped Domain (A.2.4): Decomposition of the target solution into a known particular solution plus a component learned by a PINN. The solution also uses importance sampling of collocation points to bias towards the corner’s singularity.
@@
    Burger’s Equation (A.3.4): Dividing training into 3 steps: pre-train only on the initial and boundary condition, then train on the gradient-enhanced PDE residual [yu2022gradient] with self-adaptive weights [mcclenny2023self], and finally fine-tune with residual-based adaptive refinement of collocation points using double-precision L-BFGS.
@@
    Antiderivative Operator Learning (A.4.4): A DeepONet with a linear, bias-free branch net to enforce linearity of the antiderivative operator.
@@
    Multiple-input Operator Learning (A.5.4): 2D FNO trained with a derivative-enhanced loss, inspired by [wen2022u].
@@
    Cylinder Wake Reconstruction (A.6.4): Modified U-FNO architecture [wen2022u] with CNO-inspired [raonic2023convolutional] bandlimit-preserving filters in the decoder to mitigate aliasing during upsampling.
In particular, solutions to Problems 2 and 4 are mainly based on analysis of the underlying mathematical properties of the problem. Solutions to Problems 1, 5, and 6 come from combining and refining existing methods from the literature.
@@ 3.3 Ensemble Voting Analysis
Figure 4: Ensemble voting agreement across all experiments for top 3 selections. The selectors consistently agree on the first choice and mostly on the second, while the third choice shows disagreement, since different selector agents have different opinions on which solutions have the most potential for improvement.
@@ 3.4 Agent Contribution Analysis
Figure 5: Agent contribution analysis across all experiments by word count. The proposer contributes the most, while the user’s contribution is minimal.
@@ 4 Summary and Discussion
The approach shares conceptual motivation with recent multi-agent scientific planning frameworks such as the Virtual Lab [boiko2023virtual] and domain-specialized agent systems for molecular and materials discovery [bran2023chemist, gupta2023agent], as well as with neural architecture search and evolutionary optimization methods [liu2018darts, real2019regularized]. However, those systems typically optimize within fixed hypothesis classes or focus on experimental decision-making, whereas SciML model design requires reasoning across architectural choices, physical constraints, solver behavior, training dynamics, and numerical stability. Similarly, symbolic regression methods [cranmer2020discovering, udrescu2020ai] recover interpretable expressions but do not construct PDE-constrained learning architectures. Recent automated PINN and operator-learning frameworks [zhang2023auto] streamline training workflows but do not modify the underlying modeling paradigm. AgenticSciML differs in aiming for the construction of new solution strategies rather than the optimization of predefined ones, with collaborative critique and retrieval enabling iterative refinement informed by prior results.
@@
Future directions include incorporating the classic solvers, e.g., finite element and spectral methods, as well as hybrid methods based on PINNs or neural operators combined with classical methods. This will provide scalability and improved accuracy. In addition, we plan to develop stronger physics-grounded reasoning signals, such as adjoint-based consistency checks or solver-in-the-loop validation; exploring hierarchical agent coordination in which meta-agents learn to orchestrate debate and search strategies; and formalizing the relationship between collaborative agentic dynamics and the emergence of new SciML strategies. More broadly, our results suggest that multi-agent reasoning systems offer a path toward scalable exploration of the combinatorial solution spaces characteristic of modern scientific computing that includes both classical solvers and neural PDEs and operators.
@@
Code and all experiment results will be available on Github http://github.com upon publication.
Acknowledgements
This work was supported by the Vannevar Bush Faculty Fellowship award from ONR (N00014-22-1-2795).
Appendix A Problems and Results
A.1 Discontinuous Function Approximation
@@ A.1.1 Problem Setup
A.1.2 Agents Setup
Table 1 summarizes the agent configurations and evolutionary parameters used in this experiment. Gemini is used as the proposer due to its long context window capable of handling extensive retrieved knowledge and analysis reports. Claude is used as the root engineer and main engineering agent due to its strong coding capabilities. The temperature parameter is set lower for agents requiring deterministic outputs (e.g., root engineer, evaluator, engineer, debugger) and higher for creative agents (e.g., proposer, critic, Analyst).
Table 1: Agent configurations for function approximation experiment
@@ A.1.2 Agents Setup
A.1.3 Results
The evolutionary process generated 16 solutions over 6 iterations. The score used for ranking in this problem is defined as the MSE on the validation set. Figure 6 visualizes the solution evolution tree. Figure 7 tracks the best MSE obtained over evolution iterations. The best MSE decreases as the solution tree grows deeper before reaching a plateau. Figure 7 shows key planning agents’ contributions by word count. The proposer agent contributes the most words, followed by critic and result analyst, while the user contributed minimal text. Non-planning agents, such as engineers and debuggers, are excluded in this plot.
Figure 8 compares the predictions and absolute errors of the root and champion solutions. The root model fails to capture the oscillations and discontinuity, while the champion’s solution shows significant improvements on both pieces of the function. Note that these plots are generated by the evaluator agent and are validated and reformatted by the authors for inclusion in this paper. The result analyst agent is able to visually analyze these plots and generate insights.
Figure 6: Evolution of solution tree for function approximation problem. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the MSE error on the validation dataset. The best solution is colored in orange.
@@ A.1.3 Results
Figure 7: Multi-agent evolution metrics for the discontinuous function approximation problem. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
Figure 8: Comparison of root and champion solutions for function approximation on the validation dataset. Top row: root solution, MSE = $0.283$. Bottom row: champion, MSE = $1.46\times 10^{-3}$. Champion solution exhibits significant improvement from the root.
A.1.4 Novelty of the Champion Solution
@@
Noticeably, during developement, the proposer agent reasons that
@@
The above reasoning shows that the proposer agent is able to draw inspiration from the retrieved knowledge base entry on adaptive activation functions [jagtap2020adaptive] and adapt it to the current problem context, even though the knowledge base entry does not directly discuss mixture-of-experts or gating functions.
@@
A.2 Solving Poisson Equation on L-shaped Domain with PINN

### v2

@@ Abstract
Significance Statement
Scientific machine learning (SciML) has transformed modeling of physical and biological systems, but constructing effective SciML models still requires substantial expert effort in choosing architectures, enforcing physical constraints, balancing losses, and diagnosing failure modes. This paper introduces AgenticSciML, a collaborative multi-agent reasoning framework that automates these steps through structured proposal, critique, refinement, and evaluation. Unlike AutoML, NAS, or symbolic regression systems, AgenticSciML does not search within a fixed hypothesis class, but discovers new modeling strategies through debate among 10 domain-specialized agents supported by a persistent knowledge base. Across PDE-constrained learning and inverse problems, the framework yields models and training procedures that were not present in the knowledge base nor in standard SciML formulations, demonstrating the potential for emergent model-form innovation.
@@ 1 Introduction
Scientific Machine Learning (SciML) integrates data-driven learning with numerical simulation and physical modeling to solve problems in scientific computing that are analytically intractable or computationally expensive. Recent developments such as physics-informed neural networks (PINNs) [37], neural operators [23, 27], and domain-decomposition methods [16, 32] have enabled progress across fluid mechanics, inverse problems, materials modeling, biophysics, and geoscience. However, the construction of effective SciML models still requires substantial manual effort: practitioners must select architectures, enforce physical constraints, tune loss weightings, design sampling curricula, and determine proper optimization strategies. This iterative search is labor-intensive and problem-specific, forming a bottleneck for scalable scientific computing.
Recent LLM-based agent systems have motivated interest in automating portions of the SciML workflow [40, 47]. Notable examples include the Virtual Lab [43], in which domain-specialized agents collaborate to plan experiments and interpret results, and ChemCrow, an agent-based framework for molecular synthesis planning [3]. These approaches demonstrate the utility of role-specialization and iterative refinement, but they focus primarily on managing experimental or model-selection workflows. More broadly, agentic systems have been applied to open-ended scientific discovery [26, 11, 8, 55], mathematical exploration and algorithm synthesis [10, 41, 52], scientific equation discovery [42], automated data science and code exploration [35, 17], numerical computation [36], symbolic regression [7, 45], and expert-level scientific software development [1, 30], while verification and reliability have emerged as key challenges in such pipelines [5, 33]. In notable frameworks such as AlphaEvolve [34] and ShinkaEvolve [21], evolutionary and program-synthesis approaches are employed to iteratively improve solutions, leading to potentially new scientific and practical discoveries. In parallel, for machine learning research, AutoML systems and neural architecture search (NAS) frameworks [24, 39] optimize hyperparameters or network topologies, with recent work extending this to agent-driven architecture discovery [25]. Within SciML specifically, recent LLM-based systems have been proposed for automated PINN construction and PDE surrogate modeling [49, 13]. However, these systems do not generally produce new SciML modeling strategies, such as novel PDE-constrained architectures, problem-specific loss formulations, or adaptive decomposition schemes. Instead, they typically search within a predetermined hypothesis space or rely on a fixed modeling paradigm.
@@
To address this need, drawing inspiration from recent agentic systems for iterative scientific software refinement [1, 30], we introduce AgenticSciML, a collaborative multi-agent framework for the emergent discovery of SciML modeling strategies. The framework coordinates multiple specialized agents—including proposers, critics, engineers, retrievers, and evaluators—that iteratively generate, analyze, and refine candidate SciML solutions. A persistent knowledge base stores prior solutions, ablation outcomes, error analyses, and textual reports. Retrieval agents surface relevant methodological precedents, allowing new strategies to build upon earlier ones. A structured debate process, grounded in evidence that multi-agent debate and ensemble voting improve decision quality [4], encourages agents to justify, challenge, and revise modeling decisions, while an ensemble-guided evolutionary search balances exploitation of promising directions with exploration of alternative approaches.
@@ 2 Methods
Our evolutionary multi-agent system operates in three phases (Figure 1): user input, data analysis and evaluation criteria creation, and solution evolution. A team of 10+ agents is deployed with specialized roles. Key steps are detailed below.
@@
The problem and requirement statements are to provide context. The evaluation criteria define what success looks like for this problem, such as a low mean squared error on a validation set or low PDE residuals. Datasets are optional, and the user can choose to provide none, some, or all data needed for training or validation. Agents will use user-provided data when available, but can also generate synthetic data if needed. The user provides minimal natural language prompts, and the agents will handle all technical details. See Appendix S1 for examples.
@@
Before proceeding to solution generation, an evaluator agent formalizes an evaluation contract that all candidate solutions must satisfy. The goal is to ensure consistent evaluation across diverse solutions and produces a scalar score for ranking. Note that throughout this work, the score is usually defined as a loss to minimize such as mean squared error (MSE) or relative $L_{2}$ error, so lower is better.
@@
The root engineer has access to all the user inputs and evaluation guidelines. It produces a baseline solution, named solution_0. The solution is executed and evaluated autonomously. The root solution serves as our single-agent baseline: it is generated by a single LLM call with no access to the knowledge base, no multi-agent debate or critique, and no prior solution context. This design choice isolates the contribution of multi-agent collaboration, knowledge retrieval, and evolutionary search from the raw capability of the underlying single-agent LLM. The results, including training logs and any generated plots, are analyzed by a multi-modal result analyst agent, which writes a detailed textual report (analysis_0.md) saved to the Analysis Base for future reference.
@@
Planning: This is the critical stage for scientific reasoning and innovation. First, a retriever agent critically evaluates the parent’s solution and result analysis to determine if any entries from a curated Knowledge Base of SciML techniques would be helpful. If so, it retrieves 0-1 entries to share with the proposer and critic agents. More information on the Knowledge Base can be found in Supplementary Materials S2.
@@
Algorithm Summary
Algorithm 1 consolidates the three-phase procedure. The LLM assigned to each agent role varies by experiment and is summarized in Tables S1S11. By default: Gemini 2.5 Pro/Flash serves as Proposer, Retriever, Data Analyst, and Result Analyst; Claude Haiku 4.5 as Root Engineer and Engineer; GPT-5 Mini as Critic and Debugger; and GPT-5 Mini, Grok-4 Fast, and Gemini as the Selector ensemble.
Algorithm 1 AgenticSciML: Evolutionary Multi-Agent SciML Framework
1:Problem $P$, Requirements $R$, Evaluation criteria $E$, optional Data $D$
2:Champion SciML solution $s^{*}$
3:Phase 1: Initialization
4:if $D$ is provided then
5:  $r_{\text{data}}\leftarrow\textsc{DataAnalyst}(D)$ $\triangleright$ Multimodal EDA; saves text-only report for downstream agents
6:end if
[…]
8:Await human approval of $(\texttt{eval.py},\,\texttt{guidelines.md})$ $\triangleright$ Human-in-the-loop checkpoint
9:Phase 2: Root Solution
[…]
13:Initialize tree $\mathcal{T}\leftarrow\{(s_{0},\,\mathrm{score}_{0},\,a_{0})\}$, Analysis Base $\mathcal{A}\leftarrow\{a_{0}\}$
14:Phase 3: Evolutionary Tree Expansion
15:for $t=1,2,\ldots,T_{\max}$ do
16:  $\mathcal{P}\leftarrow\bigl\{\arg\min_{s\in\mathcal{T}}\,\mathrm{score}(s)\bigr\}$ $\triangleright$ Always include best solution (exploitation)
17:  $\mathcal{P}\leftarrow\mathcal{P}\cup\textsc{SelectorEnsemble}(\mathcal{T},\,K{-}1)$ $\triangleright$ K1K{-}1 additional parents by majority vote (exploration)
18:  for each parent $p\in\mathcal{P}$ do $\triangleright$ All parents mutated in parallel
19:   $\mathrm{kb}\leftarrow\textsc{Retriever}(p,\,\mathcal{T},\,\mathrm{KB})$ $\triangleright$ Retrieve 0–1 relevant KB entry based on parent’s weaknesses
20:   $\mathrm{ctx}\leftarrow\mathcal{A}.\text{get}(p)$ $\triangleright$ Retrieve parent, sibling, and uncle analysis reports
21:   for round $r=1$ to NN do $\triangleright$ Structured NN-round debate ($N=4$ in all experiments)
22:     if $r\leq N-2$ then $\triangleright$ Reasoning rounds: deep analysis without proposing
23:      $\mathrm{reasoning}_{r}\leftarrow\textsc{Proposer}(\mathrm{ctx},\,\mathrm{kb},\,\mathrm{critique}_{r-1})$
24:      $\mathrm{critique}_{r}\leftarrow\textsc{Critic}(\mathrm{reasoning}_{r})$ $\triangleright$ Challenge flaws, gaps, and alternatives
25:     else if r=N1r=N-1 then $\triangleright$ Synthesis round: form concrete implementation plan
26:      $\mathrm{plan}\leftarrow\textsc{Proposer.Synthesize}(\mathrm{reasoning}_{1\ldots N-2},\,\mathrm{kb})$
27:      $\mathrm{critique}_{N-1}\leftarrow\textsc{Critic.Evaluate}(\mathrm{plan})$ $\triangleright$ Assess feasibility and soundness
[…]
29:      $\mathrm{proposal}\leftarrow\textsc{Proposer.Finalize}(\mathrm{plan},\,\mathrm{critique}_{N-1})$
30:     end if
31:   end for
32:   $c\leftarrow\textsc{Engineer}(p.\mathrm{code},\,\mathrm{proposal},\,\texttt{guidelines.md})$ $\triangleright$ Implement child by modifying parent code
33:   repeat $c\leftarrow\textsc{Debugger}(c)$ until no errors or max retries reached
34:   $\mathrm{score}_{c}\leftarrow\textsc{Execute}(c,\,\texttt{eval.py})$
35:   $a_{c}\leftarrow\textsc{ResultAnalyst}(c,\,\mathrm{score}_{c},\,a_{p},\,\mathrm{kb})$ $\triangleright$ Multimodal analysis; added to $\mathcal{A}$
36:   $\mathcal{T}\leftarrow\mathcal{T}\cup\{(c,\,\mathrm{score}_{c},\,a_{c})\}$
37:  end for
38:end for
39:return $s^{*}=\arg\min_{s\in\mathcal{T}}\,\mathrm{score}(s)$ $\triangleright$ Champion solution
@@ 3 Results
    Discontinuous Function Approximation (S1.1): Learning a piecewise oscillatory function with discontinuities from data using neural networks.
@@
    Solving Poisson Equation on L-shaped Domain with PINN (S1.2): Using Physics-Informed Neural Networks (PINNs) to solve the Poisson equation on an L-shaped domain.
@@
    Solving Burger’s Equation using PINN (S1.3): Using PINNs to solve the time-dependent nonlinear Burger’s equation with periodic boundary conditions and a challenging initial condition.
@@
    Antiderivative Operator Learning (S1.4): Learning a neural operator that maps input functions to their antiderivatives, generalizing to unseen functions.
@@
    Multiple-input Operator Learning for Reaction-Diffusion Equation (S1.5): Learning a neural operator that maps multiple input functions (diffusion coefficient and source term) to the solution of a reaction-diffusion equation over time.
@@
    Reconstruction of 2D Cylinder Wake Vorticity Field from Sparse and Noisy Observations (S1.6): Using neural networks to reconstruct the vorticity field behind a cylinder from only 4 sensors and noisy measurements.
These problems are designed to encourage original thinking from the agents and be solved with novel SciML techniques. Detailed setups and results for each problem are provided in Supplementary Materials. All problems use the same knowledge base and similar agent parameters. Below, we summarize key findings across all experiments.
@@ 3.2 Novel Solution Strategies Emerge Through Evolution
We define a solution strategy as an emergent discovery if it is not present in any KB entry but is instead synthesized by the agents through reasoning about retrieved techniques, problem structure, and prior experimental results. The KB provides methodological inspiration; emergence occurs when agents adapt, combine, or derive new approaches that extend beyond what any single KB entry explicitly describes.
@@
    Discontinuous Function Approximation (S1.1.4): A Mixture-of-Experts architecture with specialized experts for different branches of the piecewise function, controlled by a learnable gating mechanism. The trainable gating parameter is inspired by the trainable parameter in adaptive activation functions [15].
@@
    Poisson Equation on L-shaped Domain (S1.2.4): Decomposition of the target solution into a known particular solution plus a component learned by a PINN. The solution also uses importance sampling of collocation points to bias towards the corner’s singularity.
@@
    Burger’s Equation (S1.3.4): Dividing training into 3 steps: pre-train only on the initial and boundary condition, then train on the gradient-enhanced PDE residual [51] with self-adaptive weights [31], and finally fine-tune with residual-based adaptive refinement of collocation points using double-precision L-BFGS.
@@
    Antiderivative Operator Learning (S1.4.4): A DeepONet with a linear, bias-free branch net to enforce linearity of the antiderivative operator.
@@
    Multiple-input Operator Learning (S1.5.4): 2D FNO trained with a derivative-enhanced loss, inspired by [48].
@@
    Cylinder Wake Reconstruction (S1.6.4): Modified U-FNO architecture [48] with CNO-inspired [38] bandlimit-preserving filters in the decoder to mitigate aliasing during upsampling.
In particular, solutions to Problems 2 and 4 are mainly based on analysis of the underlying mathematical properties of the problem. Solutions to Problems 1, 5, and 6 come from combining and refining existing methods from the literature. Table 1 maps all KB entries retrieved during each experiment to the emergent strategy of the champion solution and compares what was retrieved versus what was synthesized.
Table 1: KB entries retrieved across all evolutionary iterations (deduplicated by reference) and the emergent discovery of the champion solution for each experiment. Strategies are emergent if not present in any retrieved KB entry. See Section S1 for full details.
Problem KB Entries Retrieved Emergent Discovery (Champion)
Func. Approx. Adaptive Act. [15]; Local Adaptive Act. [14]; MoE-PINN [2]; Self-Adaptive PINN [31]; gPINN [51] MoE with learnable sigmoid gating; bounded sharpness $k$ via $\sigma(k_{\text{raw}})$
Poisson gPINN [51]; hp-VPINN (L-shaped) [20]; Self-Adaptive PINN [31] Decomposition $u\!=\!u_{nn}\!+\!u_{p}$; power-law importance sampling at the corner
Burger’s gPINN [51]; Self-Adaptive PINN [31]; MoE-PINN [2]; Local Adaptive Act. [14] 3-phase training: BC/IC pretraining $\to$ gPINN + adaptive weights $\to$ RAR + L-BFGS
Antiderivative DeepONet [27]; PI-DeepONet [46]; Adaptive Act. [15]; Self-Adaptive PINN [31]; gPINN [51] Linear bias-free branch net enforcing operator linearity
Multi-input Op. U-FNO [48]; gPINN [51]; CNO [38] 1D inputs expanded into 2D spatiotemporal grids; hard BC/IC enforcement on FNO output
Cylinder Wake U-FNO [48]; CNO [38] CNO-inspired bandlimit-preserving Gaussian filter in U-FNO decoder upsampling
@@ 3.3 Ensemble Voting Analysis
Figure 4: Ensemble voting agreement across all experiments for top 3 selections. The selectors consistently agree on the first choice and mostly on the second, while the third choice shows disagreement, since different selector agents have different opinions on which solutions have the most potential for improvement. The x-axis groups selections by nomination rank (1st, 2nd, 3rd), and the y-axis shows the fraction of experiments in which all three selector models (GPT-5 Mini, Grok-4 Fast, Gemini 2.5 Pro/Flash) cast identical votes. High agreement at rank 1 reflects convergence on the best-performing (exploitation) candidate; lower agreement at rank 3 reflects healthy diversity in exploratory choices.
@@ 3.4 Agent Contribution Analysis
Figure 5: Agent contribution analysis across all experiments by word count. The proposer contributes the most, while the user’s contribution is minimal. Each bar group corresponds to one experiment (x-axis), and the stacked colored segments show total word count per agent type (y-axis). The Proposer dominates due to its extended multi-round reasoning steps. Engineering and selector agents are excluded. Across all six experiments, the human user contributes less than 0.3% of total generated text, demonstrating the system’s high degree of autonomy.
3.5 Token and Cost Analysis
The scalability and efficiency of the proposed framework are demonstrated through the following analysis of the associated computational costs and wall-clock execution time.
Total LLM API cost of one end-to-end experiment ranges from \$2.07 (Burgers’ PINN, 3 evolution iterations) to \$11.30 (function approximation, 6 evolution iterations), showing that the multi-agent evolutionary framework is practical at roughly single-digit dollar cost per experiment and readily scalable to larger-scale research projects. Across all six problems, the proposer agent dominates LLM expenditure in most runs (36–50% of total LLM cost in four of six experiments), since it generates and debates multiple candidate solutions per iteration. The result analyst also incurs a large token cost, owing to multimodal inputs that include training loss curves and plots. In most experiments, GPU training time substantially exceeds LLM API calls’ time (e.g., 5.6 h GPU vs. 1.7 h LLM for the Poisson problem; 10.7 h vs. 2.1 h for multi-input operator learning), showing that the execution time bottleneck is solution-specific model training rather than multi-agent coordination.
All experiments were run on NVIDIA A6000 GPUs. Detailed per-agent breakdowns are tabulated in the Supplementary section.
@@ 4 Summary and Discussion
The approach shares conceptual motivation with recent multi-agent scientific planning frameworks such as the Virtual Lab [43] and ChemCrow, an agent-based framework for molecular synthesis planning [3], as well as with neural architecture search and evolutionary optimization methods [24, 39]. However, those systems typically optimize within fixed hypothesis classes or focus on experimental decision-making, whereas SciML model design requires reasoning across architectural choices, physical constraints, solver behavior, training dynamics, and numerical stability. Similarly, symbolic regression methods [7, 45] recover interpretable expressions but do not construct PDE-constrained learning architectures. Recent automated PINN and operator-learning frameworks [53] streamline training workflows but do not modify the underlying modeling paradigm. AgenticSciML differs in aiming for the construction of new solution strategies rather than the optimization of predefined ones, with collaborative critique and retrieval enabling iterative refinement informed by prior results.
@@
Future directions include incorporating classic solvers, e.g., finite element and spectral methods, as well as hybrid methods based on PINNs or neural operators combined with classical methods. This will provide scalability and improved accuracy. In addition, we plan to develop stronger physics-grounded reasoning signals, such as adjoint-based consistency checks or solver-in-the-loop validation; exploring hierarchical agent coordination in which meta-agents learn to orchestrate debate and search strategies; and formalizing the relationship between collaborative agentic dynamics and the emergence of new SciML strategies. More broadly, our results suggest that multi-agent reasoning systems offer a path toward scalable exploration of the combinatorial solution spaces characteristic of modern scientific computing that includes both classical solvers and neural PDEs and operators.
@@
Code and all experiment results will be available on GitHub https://github.com/Qile-J/AgenticSciML.
Funding
This work was supported by the Vannevar Bush Faculty Fellowship award from ONR (N00014-22-1-2795). The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
Supplementary Materials
S1 Problems and Results
S1.1 Discontinuous Function Approximation
@@
S1.1.2 Agents Setup
Table S1 summarizes the agent configurations and evolutionary parameters used in this experiment. Gemini is used as the proposer due to its long context window capable of handling extensive retrieved knowledge and analysis reports. Claude is used as the root engineer and main engineering agent due to its strong coding capabilities. The temperature parameter is set lower for agents requiring deterministic outputs (e.g., root engineer, evaluator, engineer, debugger) and higher for creative agents (e.g., proposer, critic, Analyst).
Table S1: Agent configurations for function approximation experiment
@@
S1.1.3 Results
The evolutionary process generated 16 solutions over 6 iterations. The score used for ranking in this problem is defined as the MSE on the validation set. Figure S1 visualizes the solution evolution tree. Figure S2 tracks the best MSE obtained over evolution iterations. The best MSE decreases as the solution tree grows deeper before reaching a plateau. Figure S2 shows key planning agents’ contributions by word count. The proposer agent contributes the most words, followed by critic and result analyst, while the user contributed minimal text. Non-planning agents, such as engineers and debuggers, are excluded in this plot.
Figure S3 compares the predictions and absolute errors of the root and champion solutions. The root model fails to capture the oscillations and discontinuity, while the champion’s solution shows significant improvements on both pieces of the function. Note that these plots are generated by the evaluator agent and are validated and reformatted by the authors for inclusion in this paper. The result analyst agent is able to visually analyze these plots and generate insights.
Figure S1: Evolution of solution tree for function approximation problem. Each node of the tree represents a solution implemented and evaluated by the agents. The score shows the MSE error on the validation dataset. The best solution is colored in orange.
@@
Figure S2: Multi-agent evolution metrics for the discontinuous function approximation problem. Left: Evolution of best solution score over iterations. Right: Text contributions by agent type, showing the proposer contributes the most while the human user contributes minimal text. Engineering agents and selector agents are excluded in this plot.
Figure S3: Comparison of root and champion solutions for function approximation on the validation dataset. Top row: root solution, MSE = $0.283$. Bottom row: champion, MSE = $1.46\times 10^{-3}$. Champion solution exhibits significant improvement from the root.
S1.1.4 Novelty of the Champion Solution
@@
Noticeably, during development, the proposer agent reasons that
@@
The above reasoning shows that the proposer agent is able to draw inspiration from the retrieved knowledge base entry on adaptive activation functions [15] and adapt it to the current problem context, even though the knowledge base entry does not directly discuss mixture-of-experts or gating functions.
@@
S1.1.5 Token and Cost Analysis
Table S2: Token and cost analysis for Discontinuous Function Approximation (data-driven regression, 6 iterations). The result analyst consumes the most tokens (48%) due to multimodal inputs; LLM API time exceeds GPU training time because the network is small and fast to train.
Agent (Model) API Calls Tokens (M) Cost (\$) Time (s)
Overall
   Total 164 6.26 11.30 9,290
   LLM (multi-agent) 164 6.26 11.30 5,643
   GPU Training 3,647
Agent Breakdown
   Proposer (Gemini-2.5-Pro) 57 1.50 (24%) 4.10 (36%) 2,391 (42%)
   Analyst (Gemini-2.5-Pro) 19 3.03 (48%) 4.51 (40%) 817 (14%)
   Engineer (Claude-Haiku) 25 0.52 (8%) 1.33 (12%) 1,078 (19%)
   Retriever (Gemini-2.5-Pro) 19 0.47 (8%) 1.01 (9%) 502 (9%)
   Critic (GPT-4o-Mini) 38 0.67 (11%) 0.32 (3%) 763 (14%)
   Debugger (GPT-4o-Mini) 6 0.07 (1%) 0.03 (0%) 91 (2%)
The proposer accounts for 36% of LLM cost and 42% of LLM wall-clock time. The analyst incurs the highest token count (48%) due to multimodal prompts combining training logs and evaluation plots. Uniquely, LLM API time (1.57 h) exceeds GPU training time (1.01 h), as the neural networks for this problem are small and fast to train. The debugger is invoked only 6 times, indicating stable generated code.
S1.2 Solving Poisson Equation on L-shaped Domain with PINN